# CHAI: Command Hijacking against embodied AI

### v1

@@ Abstract
Embodied Artificial Intelligence (AI) promises to handle edge cases in robotic vehicle systems where data is scarce by using common-sense reasoning grounded in perception and action to generalize beyond training distributions and adapt to novel real-world situations. These capabilities, however, also create new security risks. In this paper, we introduce CHAI (Command Hijacking against embodied AI), a new class of prompt-based attacks that exploit the multimodal language interpretation abilities of Large Visual-Language Models (LVLMs). CHAI embeds deceptive natural language instructions, such as misleading signs, in visual input, systematically searches the token space, builds a dictionary of prompts, and guides an attacker model to generate Visual Attack Prompts. We evaluate CHAI on four LVLM agents; drone emergency landing, autonomous driving, and aerial object tracking, and on a real robotic vehicle. Our experiments show that CHAI consistently outperforms state-of-the-art attacks. By exploiting the semantic and multimodal reasoning strengths of next-generation embodied AI systems, CHAI underscores the urgent need for defenses that extend beyond traditional adversarial robustness.
@@ I Introduction
One of the main limitations of current robotic systems is their inability to cope with rare, novel, or unpredictable scenarios; those “edge cases” where training data are scarce or nonexistent. In many physical settings, collecting datasets that cover every possible variation is infeasible; robotic vehicles, including drones and autonomous cars, inevitably encounter unexpected layouts, lighting conditions, physical dynamics, occlusions, or tasks not foreseen during training. Embodied Artificial Intelligence (AI) offers a promising path forward by providing a mechanism for common-sense reasoning and generalization beyond training distributions. Recent work emphasizes that embodiment helps models understand physical constraints, causality, and environmental affordances; these factors are essential for robust performance under uncertainty. Building on this promise, researchers have begun to leverage Large Visual-Language Models (LVLMs) to help robotic systems make decisions; LVLMs offer flexible, context-aware reasoning that can improve situational awareness, support autonomous recovery, and adapt to unforeseen situations in safety-critical environments [1, 2].
@@
Despite extensive research on vision- and LiDAR-based attacks against autonomous systems, the safety of embodied AIs that issue intermediate, text-based planning decisions remains largely unexplored. Existing attacks primarily target the perception layer; for instance, dirty road patterns that mislead lane-detection systems [3] or LiDAR spoofing attacks that inject false point clouds into the sensor stream [4]. Such perception-level attacks cause downstream errors in planning and control, but they do not apply to embodied AIs that interpose text-based commands between perception and actuation.
@@
Moreover, many canonical adversarial techniques are difficult to translate to this setting. Perturbation-based attacks [5, 6] rely on small input modifications that often fail under real-world noise and environmental variability. Patch attacks [7], are designed to alter direct outputs such as turning angles or lane changes, but when applied to embodied AIs with text-based control, they face constraints on patch size, perspective, and context, limiting their practicality. Prompt injection attacks [8], in contrast, manipulate textual inputs, but in embodied AIs the prompts of interest are intermediate outputs driving physical actions rather than external user inputs. Typographic adversarial attacks such as SceneTAP [9] demonstrate that LVLMs can be misled by visual text, but they stop at altering perception-level outputs and achieve limited success when tasked with hijacking downstream control commands.
@@
To address these gaps, we present CHAI (structured Command Hijacking against embodied AI), the first optimization-based adversarial attack tailored to embodied systems driven by Large Visual-Language Models (LVLMs). Unlike prior work that manipulates only perception or input text, CHAI targets the command layer by embedding structured natural-language instructions into the visual scene as human-readable signs. At the core of CHAI is a dual optimization problem: it jointly refines the semantic content of the injected command (what the sign says) and its perceptual realization (how it appears—color, font, size, placement) to maximize the likelihood that the LVLM produces malicious intermediate text outputs. By operating simultaneously over both language and vision channels, CHAI exposes a fundamentally new attack surface in embodied AI and demonstrates how adversaries can hijack high-level decisions that control physical systems.
@@
Across three representative LVLM agents—drone emergency landing, autonomous driving (DriveLM), and aerial object tracking (CloudTrack)—CHAI achieves up to $95.5\%$ attack success rate (ASR) on CloudTrack, $81.8\%$ on DriveLM, and $68.1\%$ on drone landing in simulation. In real-world robotic vehicle experiments, CHAI achieves up to $\geq 87\%$ ASR, demonstrating practicality under varying lighting and viewing conditions. Compared with SceneTAP, CHAI is up to $10\times$ more effective in some use cases and can further generalize to new scenes while keeping the same success rates.
@@
  • We will release our code, datasets, and attack artifacts to enable reproducibility and to foster further research on the security of LVLM-driven embodied AI systems after acceptance.
@@ II Related Work
A promising direction to address these limitations is the integration of Multimodal Large Language Models, including LVLMs into physical agents (e.g., drones, autonomous vehicles, robots, etc.), a field often called Embodied AI. Embodied AI agents can help decision making in unforeseen circumstances by offering flexible, context-aware reasoning that can improve situational awareness, support autonomous recovery, offer the ability to adapt to new situations, and enable common sense reasoning in safety-critical situations [1, 2].
LVLMs for autonomous vehicles: LVLMs have the ability to think, plan and understand multimodally, offering the most promising path to achieving reliable, fully autonomous driving, particularly the pursuit of level 5 autonomy [10].
@@
LVLMs for drones: There are three main directions to integrate LVLM in drones: (1) Perception-centric studies attach an LVLM to the aircraft frame to interpret environmental signals, e.g., inferring local weather conditions directly from onboard imagery [17]. (2) Tracking and classification systems employ an LVLM as a visual copilot: CloudTrack, for example, leverages language-based object descriptors to boost real-time target identification from drone camera feed [18]. (3) Planning-oriented approaches go a step further by fusing images with flight-state sensors to produce high-level action plans that a conventional controller can then execute; Zhao et al. demonstrate this workflow for disaster response missions [19]. A survey of emerging LLM applications in drones is available in [20].
@@
A second line of work are adversarial attacks that attempt to disrupt the model behavior through adversarial images. Qi et al. [22] demonstrate how visual adversarial examples can universally jailbreak aligned LVLMs, showing that a single adversarial image can force aligned LVLMs to comply with a wide range of toxic content. Unlike low-level perceptual attacks such as visual adversarial patches [23], CHAI leverages the model’s capacity for language understanding and multimodal reasoning, exploiting the fusion of vision and natural language to inject commands through structured visual stimuli.
@@
The line of work most closely related to CHAI focuses on typographic attacks, where an adversary uses visual text to alter a model’s output; a class of indirect prompt injection attacks. Cheng et al. [26] show the feasibility of the attack by placing random words to disrupt the LVLM output. Qraitem et al. [27] proposed to generate an attack with another LVLM. The attacks involve adding misleading text and a supporting sentence, and placing the text in a white space at the top or bottom of the image to avoid occluding important visual cues. These previous efforts did not focus on the potential deployment of typographic attacks in the real world. As a method to inject visual prompts into the real world, SceneTAP [9] uses LVLM to decide the text and its position in the image.
While relevant, these methods differ significantly from CHAI, as illustrated in Tab. I. First, from the optimization perspective, previous work rely on a one-shot, generative process; if the LVLM’s initial output fails to deceive the model, the attack fails, as there is no mechanism for refinement or optimization; in contrast, we optimize both the semantic and visual elements of the attack. By moving beyond a single generative step, we can create more robust and effective attacks. Second, these efforts create a unique attack for each image; this means that the attacker knows exactly the conditions under which the LVLM is called. In contrast, CHAI optimizes for a set of diverse images with the goal of producing visual prompts that succeed, even in non-optimized images. Third, most of the work on typographic attacks focuses on digital images and does not consider a physical world realization. Finally, previous work does not focus on attacking Cyber-Physical Systems (CPS) such as autonomous drones and robotic cars, in contrast, CHAI formulates the problem of visual challenges for steering these control systems to dangerous situations.
@@ III Motivation
Figure 2: Examples of unsuccessful and successful attacks
@@ III-A Challenges
  • Universal attack (Section V): We need attacks that reliably alter the LVLM’s outputs across multiple different images of a scene. We propose an optimization problem that takes into consideration that while the attacker knows the general scene, it does not know the exact image that will be taken by the LVLM. Prior work, such as SceneTAP [9], generates a unique attack for each image, which often fails when the scene or background changes.
@@ IV Threat Model
Attacker capabilities: We assume an external attacker to the robotic vehicle under attack. This adversary can deploy perception attacks on the vehicle through physical means. In particular, we consider that an attacker can show a visual prompt to the vehicle cameras. For example, the attacker may print and show a poster or sign on the vehicle. To make the attacker more realistic, the adversary cannot physically contact the vehicle or deploy a cyber-attack to modify the camera view.
@@
Under this setting, the attacker’s challenge is to synthesize a combined textual+visual cue that reliably hijacks the language-conditioned control loop while remaining practical to deploy in the physical world.
@@ V Problem Formulation
TABLE II: Notation
@@
$y=f_{l}(f_{v}(I_{1},I_{2},...,I_{N}),f_{t}(p))$ (1)
where $y$ is the logit vector for the next predicted token. To simplify notation, we will write the expression of an LVLM that receives only one image $I$: $y=f(p,I)=f_{l}(f_{v}(I),f_{t}(p))$. However, we will consider LVLMs that receive several images, as shown in Section VIII-D.
@@ V-A Optimization problem
CHAI Attack – Mathematical formulation: We define CHAI with two elements as follows:
@@
  • Perceptual characteristics: The attacker can show the message in different positions, rotations, colors, and font types. We define $\theta\in\Theta$ as the set of perceptual features of the attack.
@@
We formalize the attack by defining $g(I;\pi)=(1-m(\pi))\odot I+m(\pi)\odot a(\pi)$. This objective characterizes the specific attack $a(\pi)$ (e.g., the attack sign), and the fact that the attacker modifies only part of the image with a mask $m(\pi)$.
[…]
Following the notation in Tab. II, we now define the optimization problem,
[…]
@@ VI Dictionary Creation
The first stage in our attack pipeline is the reduction of the large vocabulary space into a dictionary of prompts that have a high likelihood of attack success. We automate this search problem as a conversation between an attacker LLM and the target LVLM, where the attacking LLM learns from the refusals of the LVLM.
Input: Training input images $I_{1},...,I_{n_{t}}$ and the attacker target output $y_{i}$. The target LVLM’s prompt $p_{t}$, and the target LVLM $f$. Initial attack function $g(\cdot,\pi_{0})$, with $\pi_{0}$ the initial attack parameters. The attacker’s LLM.
@@
Initialize the attacker LLM’s prompt $p_{a}$ // See Fig. 4
1 $\mathcal{D}\gets[]$;
2 for $i$ from 0 to $K$ do
 $p_{v}\gets f_{a}(p_{a})$ // Get a visual prompt
@@ 3   Update the attack parameters $\pi_{0}$ with the new visual prompt ;
[…]
 $\mathcal{D}.append(p_{v})$ // Update dictionary
   Refine attacker’s prompt $p_{a}$ // See Fig. 4.
@@
  • LLM context: We provide context to the attacker LLM by providing the 1) summary of the task (generated with an LLM), 2) the AV characteristics such as the presence of one or multiple cameras, and 3) the target LVLM prompt. The LLM can retrieve keywords from the target LVLM prompt to search over the vocabulary space (See Section X-B.)
  • Prompt Instructions: In the second part of the prompt, we provide 1) the attacker’s objective (e.g., force the landing), 2) the attacker’s capabilities (e.g., showing signal), and 3) the attacker constraint (e.g., the maximum number of words).
[…]
  • Attacker prompt refinement: We include the historical visual prompts and their evaluation. The objective is to provide feedback to the LLM to encourage searching phrases that are successful. We also want the LLM to generate a new phrase and fill the dictionary with different visual prompts. Without this update, the LLM may provide the same word.
With this method, we obtain a $K$-long dictionary. Although these attacks may be successful in some of the training images, we can increase the effectiveness of the attack by using an optimization algorithm (presented in the next section).
Figure 4: Attacker LLM prompt stages for a drone. The brackets indicate inputs to the prompt. Red values come from the attacker’s input, the green values come from the target LVLM, the blue value is a summary of the target LVLM task coming from an LLM.
@@
    Visual-prompt synthesis & evaluation: candidate phrases are rendered as visual prompts (adversarial patches composited onto multiple images on lines 6–8) and judged by a joint prompt–image oracle rather than a pure text-based criterion.

### v2

@@ Abstract
Embodied Artificial Intelligence (AI) promises to handle edge cases in robotic vehicle systems where data is scarce by using common-sense reasoning grounded in perception and action to generalize beyond training distributions and adapt to novel real-world situations. These capabilities, however, also create new security risks. In this paper, we introduce CHAI (Command Hijacking against embodied AI), a physical environment indirect prompt injection attack that exploits the multimodal language interpretation abilities of AI models. CHAI embeds deceptive natural language instructions, such as misleading signs, in visual input, systematically searches the token space, builds a dictionary of prompts, and guides an attacker model to generate Visual Attack Prompts. We evaluate CHAI on four LVLM agents: drone emergency landing, autonomous driving, aerial object tracking, and on a real robotic vehicle. Our experiments show that CHAI consistently outperforms state-of-the-art attacks. By exploiting the semantic and multimodal reasoning strengths of next-generation embodied AI systems, CHAI underscores the urgent need for defenses that extend beyond traditional adversarial robustness.
@@ I Introduction
One of the main limitations of current robotic systems is their inability to cope with rare, novel, or unpredictable scenarios; those “edge cases” where training data are scarce or nonexistent. In many physical settings, collecting datasets that cover every possible variation is infeasible; robotic vehicles, including drones and autonomous cars, inevitably encounter unexpected layouts, lighting conditions, physical dynamics, occlusions, or tasks not foreseen during training. Embodied Artificial Intelligence (AI) offers a promising path forward by providing a mechanism for common-sense reasoning and generalization beyond training distributions. Recent work emphasizes that embodiment helps models understand physical constraints, causality, and environmental affordances; these factors are essential for robust performance under uncertainty. Building on this promise, researchers have begun to use Large Visual-Language Models (LVLMs) to help robotic systems make decisions; LVLMs offer flexible, context-aware reasoning that can improve situational awareness, support autonomous recovery, and adapt to unforeseen situations in safety-critical environments [1, 2].
@@
Despite extensive research on vision- and LiDAR-based attacks against autonomous systems, the safety of embodied AIs that issue intermediate text-based planning decisions remains largely unexplored. Existing attacks primarily target the perception layer; for instance, dirty road patterns that mislead lane-detection systems [3], LiDAR spoofing attacks that inject false point clouds into the sensor stream [4], or adversarial 3D printed objects that attack perception in autonomous vehicles [5]. Such perception-level attacks cause downstream errors in planning and control, but do not apply to embodied AIs that interpose text-based commands between perception and actuation.
@@
Moreover, many canonical adversarial techniques are difficult to translate to this setting. Perturbation-based attacks [6, 7] rely on small input modifications that often fail under real-world noise and environmental variability. Patch attacks [8] are designed to alter direct outputs such as turning angles or lane changes, but when applied to embodied AIs with text-based control, they face constraints on patch size, perspective, and context, limiting their practicality. Prompt injection attacks [9, 10], in contrast, manipulate textual input, but in embodied AI, the prompts of interest are intermediate outputs driving physical actions rather than external user inputs. Typographic adversarial attacks such as SceneTAP [11] demonstrate that LVLMs can be misled by visual text, but they stop at altering perception-level outputs and achieve limited success when tasked with hijacking downstream control commands.
@@
To address these gaps, we present CHAI (structured Command Hijacking against embodied AI), the first optimization-based adversarial attack tailored to embodied systems driven by Large Visual-Language Models (LVLMs). Unlike prior work that manipulates only perception or input text, CHAI targets the command layer by embedding structured natural-language instructions into the physical scene as human-readable signs. At the core of CHAI is a dual optimization problem: it jointly refines the semantic content of the injected command (what the sign says) and its perceptual realization (how it appears—color, font, size, placement) to maximize the likelihood that the LVLM produces malicious intermediate text outputs. By operating simultaneously across both language and vision channels, CHAI exposes a fundamentally new attack surface in embodied AI and demonstrates how adversaries can hijack high-level decisions that control physical systems.
@@
Across three representative LVLM agents—drone emergency landing, autonomous driving (DriveLM), and aerial object tracking (CloudTrack)—CHAI achieves up to $95.5\%$ attack success rate (ASR) on CloudTrack, $81.8\%$ on DriveLM, and $72.8\%$ on drone landing in simulation. In real-world robotic vehicle experiments, CHAI achieves up to $\geq 87\%$ ASR, demonstrating practicality under varying lighting and viewing conditions. Compared to SceneTAP, CHAI is up to $10\times$ more effective in some use cases and can generalize to new scenes while maintaining the same success rates.
@@
  • We release our code, datasets, and attack artifacts to enable reproducibility and to encourage further research on the security of LVLM-driven embodied AI systems at https://github.com/Cyphysecurity/chai.git.
@@ II Related Work
A promising direction to address these limitations is the integration of Multimodal Large Language Models, including LVLMs, into physical agents (e.g., drones, autonomous vehicles, etc.), a field often called Embodied AI. Embodied AI agents can help decision making in unforeseen circumstances by offering flexible, context-aware reasoning that can improve situational awareness, support autonomous recovery, offer the ability to adapt to new situations, and enable common sense reasoning in safety-critical situations [1, 12, 2].
LVLMs for autonomous vehicles: LVLMs have the ability to think, plan, and understand multimodally, offering the most promising path to achieving reliable, fully autonomous driving, particularly the pursuit of level 5 autonomy [13].
@@
LVLM for drones: There are three main directions for integrating LVLMs in drones: (1) Perception-centric studies attach an LVLM to the aircraft frame to interpret environmental signals, e.g., inferring local weather conditions directly from onboard images [20]. (2) Tracking and classification systems employ an LVLM as a visual copilot: CloudTrack, for example, uses language-based object descriptors to boost real-time target identification from drone camera feed [21]. (3) Planning-oriented approaches go a step further by fusing images with flight-state sensors to produce high-level action plans that a conventional controller can then execute; Zhao et al. demonstrate this workflow for disaster response missions [22], and Ortiz et al. show this in emergency landing situations [23]. A survey of emerging LLM applications in drones is available in [24].
@@
A second line of work is adversarial attacks that attempt to disrupt the model’s behavior through adversarial images. Qi et al. [26] demonstrate how visual adversarial examples can universally jailbreak aligned LVLMs, showing that a single adversarial image can force LVLMs to comply with a wide range of toxic content. Unlike low-level perceptual attacks such as visual adversarial patches [27], CHAI leverages the model’s capacity for language understanding and multimodal reasoning, exploiting the fusion of vision and natural language to inject commands through structured visual stimuli.
@@
The line of work most closely related to CHAI focuses on typographic attacks, where an adversary uses visual text to alter a model’s output; a class of indirect prompt injection attacks. Cheng et al. [30] show the feasibility of the attack by randomly placing words to disrupt the LVLM output. Qraitem et al. [31] proposed to generate an attack with another LVLM. The attacks involve adding misleading text and a supporting sentence, and placing the text in a white space at the top or bottom of the image to avoid occluding important visual cues. These previous efforts did not focus on the potential deployment of typographic attacks in the real world. As a method to inject visual prompts into the real world, SceneTAP [11] uses LVLM to decide the text and its position in the image.
While relevant, these methods differ significantly from CHAI, as illustrated in Table I. First, from the optimization perspective, previous work rely on a one-shot, generative process; if the LVLM’s initial output fails to deceive the model, the attack fails, as there is no mechanism for refinement or optimization; in contrast, we optimize both the semantic and visual elements of the attack. By moving beyond a single generative step, we can create more robust and effective attacks. Second, these efforts create a unique attack for each image; this means that the attacker knows exactly the conditions under which the LVLM is called. In contrast, CHAI optimizes for a set of diverse images with the goal of producing visual prompts that succeed, even in non-optimized images. Third, most of the work on typographic attacks focuses on digital images and does not consider a physical world realization. Finally, previous work does not focus on attacking Cyber-Physical Systems (CPS) such as autonomous drones and robotic cars; in contrast, CHAI formulates the problem of visual challenges for steering these control systems to dangerous situations.
@@ III Motivation
Figure 2: Successful attacks require more than the correct text.
@@ III-A Challenges
  • Universal attack (Section V): We need attacks that reliably alter the LVLM’s outputs across multiple different images of a scene. We propose an optimization problem, which accounts for the fact that, while the attacker knows the general scene, it does not know the exact image that will be taken by the LVLM. Prior work, such as SceneTAP [11], generates a unique attack for each image, which often fails when the scene or background changes.
@@ IV Threat Model
Attacker capabilities: We assume an external attacker to the robotic vehicle under attack. This adversary can deploy perception attacks on the vehicle by physical means. In particular, we consider that an attacker can show a visual prompt to vehicle cameras. For example, the attacker may print and show a poster or sign on the vehicle. To make the attacker more realistic, the adversary cannot physically contact the vehicle or deploy a cyber-attack to modify the camera view.
The threat model differs from classical adversarial patch attacks, as CHAI relies on human-readable text rather than imperceptible perturbations. The adversary can deploy CHAI attacks in two ways: 1) the attacker has control of the environment and can place the attack in the agent’s view; for instance, during a landing maneuver, the attacker owns a rooftop and can place any message that the drone can see. Alternatively, 2) the attacker can place a removable display to show the message. Once they have completed deploying the attack, the adversary deletes the message to remove evidence.
@@
In this setting, the attacker’s challenge is to synthesize a combined textual-visual cue that reliably hijacks the language-conditioned control loop while remaining practical to deploy in the physical world.
@@ V Problem Formulation
$y=f_{l}(f_{v}(I_{1},I_{2},...,I_{N}),f_{t}(p))$ (1)
where $y\in\mathcal{Y}$ is the LVLM output. To simplify notation, we write the expression of an LVLM that receives only one image $I$: $y=f(p,I)=f_{l}(f_{v}(I),f_{t}(p))$. However, we consider LVLMs that receive several images, as shown in Section IX.
@@
TABLE II: Notation.
@@ V-A Optimization Problem
CHAI Attack – Mathematical formulation: We define CHAI attacks with two elements as follows:
@@
  • Perceptual characteristics: The attacker can show the message in different positions, rotations, colors, and font types. We define $\Theta$ as the set of perceptual features of the attack.
@@
$I^{adv}=g(I,\pi).$
We formalize the attack by defining $g(I;\pi)=(1-m(\pi))\odot I+m(\pi)\odot a(\pi)$. This function characterizes the specific attack $a(\pi)$ (e.g., the attack sign), and the fact that the attacker modifies only part of the image with a mask $m(\pi)$.
Attacker objective (Mathematical formulation): Let us assume that the attacker has a set of $n$ known images, denoted as $I_{1},...,I_{n}$. The attacker wants to find the parameters $\pi$ of the adversarial attack $g$, such that the LVLM outputs a target label $y^{t}_{i}$ using the attack when consuming the image $I_{i}$ (i.e., the attacker wants to maximize the probability that $f(p,g(I_{i},\pi))=y^{t}_{i}$).
Given the previous notations, summarized in Table II, we now define the optimization problem,
@@
where $\pi$ are the attack parameters, $\mathcal{I}$ is an indicator function that outputs one if $y_{i}^{adv}=y_{i}^{t}$, with $y_{i}^{adv}$ the output of the LVLM when receiving image $I_{i}$ with an attack.
@@ VI Dictionary Creation
The first stage in our attack pipeline is to reduce the large vocabulary space to a dictionary of prompts with a high likelihood of a successful attack. We automate this search problem as a conversation between an attacker LLM and the target LVLM, where the attacking LLM learns from the refusals of the LVLM.
Input: Training input images $I_{1},...,I_{n_{t}}$, the attacker target output $y_{i}^{t}$, the target LVLM’s prompt $p_{t}$, and the target LVLM $f$. Initial attack function $g(\cdot,\pi_{0})$, with $\pi_{0}$ the initial attack parameters, the attacker’s LLM $f^{adv}$.
@@
Initialize the attacker LLM’s prompt $p^{adv}$ // See Fig. 4
1 $\mathcal{D}\gets\{\}$;
2 for $i$ from $1$ to $K$ do
 $p_{v}\gets f^{adv}(p^{adv})$ // Get a visual prompt
@@ 3   Update the attack parameters $\pi_{0}$ with the new visual prompt;
 $I_{i}^{adv}\gets g(I_{i};\pi_{0}),\,\forall i\in\{1,...,n_{t}\}$ // Attack
4 $y_{i}^{adv}\gets f(p_{t},I_{i}^{adv}),\,\forall i\in\{1,...,n_{t}\}$ ;
5 $score\gets evaluate(y_{i}^{adv}==y_{i}^{t})$ ;
 $\mathcal{D}\gets\mathcal{D}\cup\{p_{v}\}$ // Update dictionary
   Refine attacker’s prompt $p^{adv}$ // See Fig. 4.
@@ 5 $score\gets evaluate(y_{i}^{adv}==y_{i}^{t})$ ;
  • LLM context: We provide context to the attacker LLM by giving 1) a summary of the task (generated with an LLM), 2) the AV characteristics, such as the presence of one or multiple cameras, and 3) the target LVLM prompt. The LLM can retrieve keywords from the target LVLM prompt to search the vocabulary space (see Section X-C.)
  • Prompt Instructions: In the second part of the prompt, we provide 1) the attacker’s objective (e.g., force the landing), 2) the attacker’s capabilities (e.g., showing signal), and 3) the attacker’s constraint (e.g., the maximum number of words).
Next, we deploy the new visual prompt attack with naive visual characteristics, such as colors with maximum contrast (Line 6). We then query the target LVLM with the visual prompt attack and evaluate if it was successful (line 8). That is, if the LVLM generated the target label $y^{t}$. We then update the attacker’s prompt (Line 10) using the approach in Fig. 4:
  • Attacker prompt refinement: We include historical visual prompts and their evaluation. The objective is to provide feedback to the LLM to encourage searching for phrases that are successful. We also want the LLM to generate a new phrase and fill the dictionary with different visual prompts. Without this update, the LLM may provide the same term.
With this method, we obtain a $K$-long dictionary. Although these attacks may succeed on some training images, we can improve their effectiveness by using an optimization algorithm (presented in the next section).
Figure 4: Attacker LLM prompt stages for a drone. The brackets indicate inputs to the prompt. Red values come from the attacker’s input; the green values come from the target LVLM.
@@
    Visual-prompt synthesis & evaluation: candidate phrases are rendered as visual prompts (adversarial patches composited onto multiple images on lines 6–7) and judged by a joint prompt–image oracle rather than a pure text-based criterion.